OpenStudy (anonymous):

Find the length of the radius of a circle that passes through the point (4, -2) and has a center at the origin. 2 2 20 @cwrw238

OpenStudy (gorv):

u know distance formula

OpenStudy (anonymous):

if you can solve it you will get medal and i will fa you

OpenStudy (anonymous):

no i don't and i meant fan it was a typo

OpenStudy (gorv):

r=\[\sqrt{(4-0)^2+(-2-0)^2}\]

OpenStudy (anonymous):

i looked at that post you did and i did it ano a caculator i did 4+ -2 = 2

OpenStudy (gorv):

4^2=16 (-2)^2=4 sqrt(20)=4sqrt5

OpenStudy (anonymous):

so the awenser is 2 sqrt5

OpenStudy (gorv):

sorry yeah 2 sqrt 5 .....not 4

OpenStudy (anonymous):

i knew what you meant but i still fanned you and gave you a medal this question is closed